Questions 11, 12  and  13  are  based  on  the  following  information: If a marginal revenue function of a rm is given as

MR = 10Q2 + 6Q − 3,

 where Q is the number of units sold.


11

Find an expression for the total revenue function (TR).

  1. 10/3 Q3 + 3Q2 - 3Q + C
  2. 5/3 Q3 + 6Q2 - 3Q + C
  3. Q3 + 3Q2 - 3Q + C
  4. 3/5 Q3 + 3Q2 - 3Q + C


12

Find the total revenue function if it is given that TR = 0 when Q = 0.

  1. 10/3 Q3 + 3Q2 - 3Q
  2. 5/3 Q3 + 6Q2 - 3Q + 1
  3. Q3 + 3Q2 - 3Q + 3
  4. 3/5 Q3 + 3Q2 - 3Q

   13

Calculate the firm's total revenue when 5 units are sold.


[1      310,36

[2      200,00

[3      476,67

[4      532,67


1
Expert's answer
2021-07-07T05:11:31-0400

Question 11

Integrating MR

"\int MR =\int( 10Q^2 + 6Q − 3)= TR"

"TR = \frac{10}{3}Q^3+3Q^2-3Q+C"

Option 1 is correct


Question 12

"TR = \frac{10}{3}Q^3+3Q^2-3Q+C"

"0= \frac{10}{3}*0^3+3*0^2-3*0+C \implies C=0"

"TR = \frac{10}{3}Q^3+3Q^2-3Q"

Option 1 is correct


Question 13

"TR = \frac{10}{3}Q^3+3Q^2-3Q"

"TR = \frac{10}{3}*5^3+3*5^2-3*5"

"TR = 476.66"

Option 3 is correct